Mike and Sarah buy books at a sale. Mike buys 3 hardcover books and 4 paperback books for $6.50. Sarah buys 2 hardcover books and 6 paperback books for $6.00. What is the cost, in dollars, of each hardcover book?

Answer by checkley79(3341)

3H+4P=6.50 MULTIPLY BY -1.5 & ADD.
2H+6P=6.00
-4.5H-6P=-9.75
-----------------------
-2.5H=-3.75
H=-3.75/-2.5
H=$1.50 COST OF EACH HARDCOVER BOOK.
3*1.50+4P=6.50
4.50+4P=6.50
4P=6.50-4.50
4P=2.00
P=2.00/4
P=$.50 COST OF EACH PAPERBACK BOOK.
PROOF:
2*1.50+6*.50=6.00
3.00+3.00=6.00
6.00=6.00